OpenStudy (anonymous):

The ratio of corresponding sides of similar triangles is four colon nine. If the longest side of the smaller triangle is thirty-six mm, what is the longest side of the larger triangle? 4 mm 16 mm 72 mm 81 mm

OpenStudy (anonymous):

36/x = 4/9 x = 81 mm

OpenStudy (anonymous):

81 is right

OpenStudy (anonymous):

Thank you...

OpenStudy (ash2326):

Let the longest side of smaller triangle be x and the longest side of larger triangle be y we are given the side ratio as 4/9 side of small triangle by side of larger triangle is 4/9 (it's not the reverse because 4/9 <1) \[\frac xy=\frac 49\] we have x as 36 so \[\frac{36}y=\frac 49\] I think you can find now!!!

OpenStudy (ash2326):

@prettyandpaid123 do you understand this?